A firm is considering the acquisition of a new machine. The base price is $85,000 and it would cost $15,000 to install. The machine is MACRS 3 year class property and it will be sold after 3 years for $17,000. The machine would also require an increase in net working capital of $10,000. The machine is expected to increase before tax revenues by $40,000 per year. This firm is in a 34% marginal tax bracket. MACRS 3 year factors are 33%, 45%, 15%, and 7% for years 1 through 4 respectively. What is the year 2 operating cash flow.

a 110,000

b 26400

c 15300

d 41700
Answer #1

rate positively ..

Computation of year 2 operating cash flow
i Revenue = 40000
ii Depreciation = 45000
(85000+15000)*45%
iii=i-ii Profit before tax = -5000
iv=iii*40% Tax @ 34% -1700
v=iii-iv Profit after tax = -3300
vi=v+ii Operating cash flow 41700
Correct answer is option d) 41700